# Characterizing and modeling harms from interactions with design patterns in AI interfaces

### v1

@@ 1. Introduction
Decades-long trends have shown that the success and popularity of a technology depend on the usability, accessibility, and user-friendliness of its interface (Godwin-Jones, 2008; Knutson, 1997). This trend is also evident in the development of new applications with artificial intelligence (AI) systems; scholars have suggested that TikTok’s broad popularity should be attributed not only to its recommendation algorithm but also to its vertical video design, which, among other things, makes bad recommendations less noticeable (Narayanan, 2022; Striano, 2023). Analyses of ChatGPT’s virality regularly point to its minimalistic and unobtrusive user interface as being a key driver of its success (CONVERSATION, 2023; Ramlochan, 2023).
@@ 2.3. Conceptualizing Risks & Harms in Sociotechnical Systems
Research on the negative impact of AI systems has expanded in response to their growing deployment. This research broadly conceptualizes harm as the negative outcomes of “entangled dynamics between design decisions, norms, and power”. (Shelby et al., 2023, p. 2) Researchers have focused on both identifying and categorizing existing harms as well as anticipatory risks of various harms, including representational, economic, and social harms (Shelby et al., 2023; Weidinger et al., 2022). Researchers have also developed measurement tools such as impact assessments, algorithmic audits, and various model evaluation approaches (Raji et al., 2020; Metcalf et al., 2021; Solaiman et al., 2023; buccinca2023aha). In this research, the assessment of AI systems’ negative impact involves creating indicators (such as error rates or bias measurements) that represent potential harms, and that can be used can be used to affect certain regulatory or technical decisions (Metcalf et al., 2021).
Research on AI impact and harm increasingly draws on HCI methods to assess how human behaviors and interactions with AI systems complicate safe and responsible AI practices (Liao and Sundar, 2022; Araujo et al., 2020). In HCI research, a significant area of study is the study of ‘dark patterns,’ a term originally coined by Brignull, which directly examines the harms inflicted by digital interface designs (brignull2015dark). In this work, harm is understood as user-centric, interaction-focused, and conceptualized more narrowly as direct impact on individuals. Early dark patterns studies focused on privacy and financial harms caused by features such as intrusive pop-ups in cookie banners (Soe et al., 2020; Gray et al., 2021b), deceptive pricing layouts in online shopping websites (Mathur et al., 2019; Voigt et al., 2021), and compulsive reward systems in video games (Zagal et al., 2013; Goodstein, 2021). More recent studies have broadened to other domains and different types of harm; some researchers investigate “attention-capture” harms, showing that dark patterns that lead to addictive consumption are prevalent on social media platforms (Monge Roffarello et al., 2023; Mildner et al., 2023). Others examine dark patterns across internet-of-things devices, highlighting many instances of privacy harms (Kowalczyk et al., 2023).In our work, we extend these studies to analyze interface designs in AI applications with a focus on the harms they pose to individual welfare and the feedback loops of human-AI behavior that could make them distinct.
@@ 3.2. Results
Multiple reviewed studies discuss design patterns in AI interfaces that have been previously categorized as dark patterns. These studies explicitly identify design features as “dark patterns” or use the names of recognized dark pattern strategies, as articulated by Gray et al., such as obstruction or sneaking (Gray et al., 2018). For example, Zhang et al. explore user privacy concerns with ChatGPT finding that obstruction is employed to dissuade users from altering their data usage settings (Zhang et al., 2023); they show that to prevent chat data from being used in training future models, users must either lose access to the chat history feature or navigate a hidden request within an FAQ page to retain it—a task none of their study participants could accomplish. Alberts and Van Kleek examine phone notifications from social agents, showing that nagging, a dark pattern involving persistent and intrusive prompts, is used to force users into “social situations” with these agents, contributing to anthropomorphic perceptions (Alberts and Van Kleek, 2023). And, in interfaces with recommendation systems, several dark patterns were described, including those concerning time unawareness, such as hiding time on the screen, and extreme countdowns, such as autoplay timers (Chaudhary et al., 2022). Many of these traditional dark patterns in human-AI interfaces seem to serve data extraction purposes, either through increasing engagement or obfuscating privacy controls (Lacey and Caudwell, 2019; Zhang et al., 2023; leschanowsky2023privacy).
@@ 3.2.2. Anthropomorphism
In human-AI interaction, anthropomorphism can be defined as attributing human traits, emotions, or intentions to AI systems (Salles et al., 2020). The anthropomorphism of AI systems by users is generally described as the outcome of anthropomorphic cues interacting with existing user mental models of these systems (Papagni and Koeszegi, 2021; Hwang and Won, 2022). Anthropomorphic cues manifest in various forms. First, cues manifest in the appearance or visualization of systems. For example, humanoid representations may lead to attributions of sentience and an overassignment of agency (Hwang and Won, 2022; Papagni and Koeszegi, 2021). In contrast, Lacey and Caudwell identify “cuteness as a dark pattern” and how it can result in attributing powerlessness to systems (Lacey and Caudwell, 2019). Second, cues may be embedded in interfaces in the form of “priming statements” that humanize AI agents. Pataranutaporn et al. find that priming statements portraying an AI system’s motive as good, evil, or neutral, shape user mental models and interactions, even when the AI model itself remains unchanged (Pataranutaporn et al., 2023). The effect they observe is even more pronounced with generative models. Finally, cues can be present in the natural language output of AI systems, like conversational or language models, where they may create ambiguity between social and transactional interactions. This may occur by incorporating language that references social relations, feelings, and emotions, effectively blurring the distinction between a user interacting with an AI system and with a human (Zhan et al., 2023; Clark et al., 2019; abercrombie2023mirages).
@@ 3.2.4. Seamless design and lack of friction
Typically, user interfaces exhibit frictionlessness through seamless designs. For example, in online platforms, such designs include features like autoplay, infinite scrolling, and the absence of context around presented content (Masrani et al., 2023; Chaudhary et al., 2022; Brady et al., 2023; Kender and Frauenberger, 2022; albuquerque2023social). While seamless designs can streamline user experience and positively enhance it, they may also inadvertently obscure critical reflection and informed decision-making. In the context of AI systems, the literature on recommendation systems in online platforms frequently cites the lack of friction caused by seamless designs on these platforms as a source of mismatch between user preferences and user actions. The concept of frictionlessness is often discussed in relation to ‘system 1’ and ‘system 2’ behaviors, terms initially coined by Kahneman (Kahneman, 2011). Frictionlessness tends to encourage system 1 behaviors, which research in psychology and cognition describes as automatic and reactive (i.e., thinking fast), as opposed to system 2 behaviors, which are more intentional and reflective (i.e., thinking slow), aligning more closely with people’s true preferences (Kahneman, 2011; Masrani et al., 2023; Sharevski et al., 2021).
@@ 4.1.1. Generating hypotheses for iterative and empirical assessments
Current impact assessment frameworks prioritize the identification, evaluation, and mitigation of harms associated with the use of AI systems (Metcalf et al., 2021; buccinca2023aha). DECAI aims to facilitate the early stages of these assessments, specifically targeting interface designs. It can assist practitioners in systematically identifying particular design features for detailed examination and provides a structured approach for formulating focused, testable hypotheses suitable for empirical study. DECAI also implicitly considers the diverse conditions of users, such as their technical proficiency and emotional states, in that process ensuring that hypothesis generation takes various user experiences into account.
@@ 4.1.2. Motivating the utility of control systems theory
We draw on control systems theory due to both its pragmatic and conceptual relevance. First, this field of research emphasizes the dynamic study and control of multi-component systems that are inherently variable, mirroring the sociotechnicality and adaptability of human-AI systems (Anand, 2013). Second, it highlights control and moderation within systems—two themes conceptually connected to autonomy-undermining designs and research on AI harm and deceptive designs more broadly (Lyshevski, 2001; Boyd, 2022; Birhane et al., 2022; Gray et al., 2021a). Increasingly, HCI research has grappled with the challenges of designing user experiences for AI systems particularly due to (1) the uncertainty of AI systems’ capabilities and (2) the complexity of AI systems’ outputs, which may range from simple to adaptively complex (Yang et al., 2020; dove2017ux; yang2016planning). This context of variability and complexity underscores the relevance of DECAI’s approach, which integrates control systems theory to respond to the dynamic nature of sociotechnical systems and their often overlooked feedback loops.
@@ 6. Discussion
Regulatory opportunities. Audits of dark patterns have gained notable regulatory traction, being codified into both EU and US law (European Parliament and the Council of the European Union, 2023b, a; European Commission, 2023a). We believe that our work can aid further research and evidence collection building on this momentum in the AI context (king2021regulating). Currently, AI-focused legislation such as the EU AI Act only indirectly addresses interface designs through regulations on transparency, human oversight, and AI-driven manipulation (European Commission, 2023b). However, interface designs may allow for the circumvention of such regulation, for example, by allowing for the strategic placement of transparency disclosures to diminish their visibility. Our results also point to how harmful designs can be used to further data collection, a practice central to today’s AI industry (Kak and West, 2023). In that way, interface designs can be pivotal in shaping both current and future regulations aimed at safe and responsible AI practices.

### v2

@@ 1. Introduction
Decades-long trends have shown that the success and popularity of a technology depend on the usability, accessibility, and user-friendliness of its interface (Godwin-Jones, 2008; Knutson, 1997). This trend is also evident in the development of new applications with artificial intelligence (AI) systems; scholars have suggested that TikTok’s broad popularity should be attributed not only to its recommendation algorithm but also to its vertical video design, which, among other things, makes bad recommendations less noticeable (Narayanan, 2022; Striano, 2023). Analyses of ChatGPT’s virality regularly point to its minimalistic and unobtrusive user interface as being a key driver of its success (Conversation), 2023; Ramlochan, 2023).
@@ 2.3. Conceptualizing Risks & Harms in Sociotechnical Systems
Research on the negative impact of AI systems has expanded in response to their growing deployment. This research broadly conceptualizes harm as the negative outcomes of “entangled dynamics between design decisions, norms, and power”. (Shelby et al., 2023, p. 2) Researchers have focused on both identifying and categorizing existing harms as well as anticipatory risks of various harms, including representational, economic, and social harms (Shelby et al., 2023; Weidinger et al., 2022). Researchers have also developed measurement tools such as impact assessments, algorithmic audits, and various model evaluation approaches (Raji et al., 2020; Metcalf et al., 2021; Solaiman et al., 2023; Buçinca et al., 2023). In this research, the assessment of AI systems’ negative impact involves creating indicators (such as error rates or bias measurements) that represent potential harms, and that can be used can be used to affect certain regulatory or technical decisions (Metcalf et al., 2021).
Research on AI impact and harm increasingly draws on HCI methods to assess how human behaviors and interactions with AI systems complicate safe and responsible AI practices (Liao and Sundar, 2022; Araujo et al., 2020). In HCI research, a significant area of study is the study of ‘dark patterns,’ a term originally coined by Brignull, which directly examines the harms inflicted by digital interface designs (Brignull et al., 2015). In this work, harm is understood as user-centric, interaction-focused, and conceptualized more narrowly as direct impact on individuals. Early dark patterns studies focused on privacy and financial harms caused by features such as intrusive pop-ups in cookie banners (Soe et al., 2020; Gray et al., 2021b), deceptive pricing layouts in online shopping websites (Mathur et al., 2019; Voigt et al., 2021), and compulsive reward systems in video games (Zagal et al., 2013; Goodstein, 2021). More recent studies have broadened to other domains and different types of harm; some researchers investigate “attention-capture” harms, showing that dark patterns that lead to addictive consumption are prevalent on social media platforms (Monge Roffarello et al., 2023; Mildner et al., 2023). Others examine dark patterns across internet-of-things devices, highlighting many instances of privacy harms (Kowalczyk et al., 2023).In our work, we extend these studies to analyze interface designs in AI applications with a focus on the harms they pose to individual welfare and the feedback loops of human-AI behavior that could make them distinct.
@@ 3.2. Results
Multiple reviewed studies discuss design patterns in AI interfaces that have been previously categorized as dark patterns. These studies explicitly identify design features as “dark patterns” or use the names of recognized dark pattern strategies, as articulated by Gray et al., such as obstruction or sneaking (Gray et al., 2018). For example, Zhang et al. explore user privacy concerns with ChatGPT finding that obstruction is employed to dissuade users from altering their data usage settings (Zhang et al., 2023); they show that to prevent chat data from being used in training future models, users must either lose access to the chat history feature or navigate a hidden request within an FAQ page to retain it—a task none of their study participants could accomplish. Alberts and Van Kleek examine phone notifications from social agents, showing that nagging, a dark pattern involving persistent and intrusive prompts, is used to force users into “social situations” with these agents, contributing to anthropomorphic perceptions (Alberts and Van Kleek, 2023). And, in interfaces with recommendation systems, several dark patterns were described, including those concerning time unawareness, such as hiding time on the screen, and extreme countdowns, such as autoplay timers (Chaudhary et al., 2022). Many of these traditional dark patterns in human-AI interfaces seem to serve data extraction purposes, either through increasing engagement or obfuscating privacy controls (Lacey and Caudwell, 2019; Zhang et al., 2023; Leschanowsky et al., 2023).
@@ 3.2.2. Anthropomorphism
In human-AI interaction, anthropomorphism can be defined as attributing human traits, emotions, or intentions to AI systems (Salles et al., 2020). The anthropomorphism of AI systems by users is generally described as the outcome of anthropomorphic cues interacting with existing user mental models of these systems (Papagni and Koeszegi, 2021; Hwang and Won, 2022). Anthropomorphic cues manifest in various forms. First, cues manifest in the appearance or visualization of systems. For example, humanoid representations may lead to attributions of sentience and an overassignment of agency (Hwang and Won, 2022; Papagni and Koeszegi, 2021). In contrast, Lacey and Caudwell identify “cuteness as a dark pattern” and how it can result in attributing powerlessness to systems (Lacey and Caudwell, 2019). Second, cues may be embedded in interfaces in the form of “priming statements” that humanize AI agents. Pataranutaporn et al. find that priming statements portraying an AI system’s motive as good, evil, or neutral, shape user mental models and interactions, even when the AI model itself remains unchanged (Pataranutaporn et al., 2023). The effect they observe is even more pronounced with generative models. Finally, cues can be present in the natural language output of AI systems, like conversational or language models, where they may create ambiguity between social and transactional interactions. This may occur by incorporating language that references social relations, feelings, and emotions, effectively blurring the distinction between a user interacting with an AI system and with a human (Zhan et al., 2023; Clark et al., 2019; Abercrombie et al., 2023).
@@ 3.2.4. Seamless design and lack of friction
Typically, user interfaces exhibit frictionlessness through seamless designs. For example, in online platforms, such designs include features like autoplay, infinite scrolling, and the absence of context around presented content (Masrani et al., 2023; Chaudhary et al., 2022; Brady et al., 2023; Kender and Frauenberger, 2022; Albuquerque et al., 2023). While seamless designs can streamline user experience and positively enhance it, they may also inadvertently obscure critical reflection and informed decision-making. In the context of AI systems, the literature on recommendation systems in online platforms frequently cites the lack of friction caused by seamless designs on these platforms as a source of mismatch between user preferences and user actions. The concept of frictionlessness is often discussed in relation to ‘system 1’ and ‘system 2’ behaviors, terms initially coined by Kahneman (Kahneman, 2011). Frictionlessness tends to encourage system 1 behaviors, which research in psychology and cognition describes as automatic and reactive (i.e., thinking fast), as opposed to system 2 behaviors, which are more intentional and reflective (i.e., thinking slow), aligning more closely with people’s true preferences (Kahneman, 2011; Masrani et al., 2023; Sharevski et al., 2021).
@@ 4.1.1. Generating hypotheses for iterative and empirical assessments
Current impact assessment frameworks prioritize the identification, evaluation, and mitigation of harms associated with the use of AI systems (Metcalf et al., 2021; Buçinca et al., 2023). DECAI aims to facilitate the early stages of these assessments, specifically targeting interface designs. It can assist practitioners in systematically identifying particular design features for detailed examination and provides a structured approach for formulating focused, testable hypotheses suitable for empirical study. DECAI also implicitly considers the diverse conditions of users, such as their technical proficiency and emotional states, in that process ensuring that hypothesis generation takes various user experiences into account.
@@ 4.1.2. Motivating the utility of control systems theory
We draw on control systems theory due to both its pragmatic and conceptual relevance. First, this field of research emphasizes the dynamic study and control of multi-component systems that are inherently variable, mirroring the sociotechnicality and adaptability of human-AI systems (Anand, 2013). Second, it highlights control and moderation within systems—two themes conceptually connected to autonomy-undermining designs and research on AI harm and deceptive designs more broadly (Lyshevski, 2001; Boyd, 2022; Birhane et al., 2022; Gray et al., 2021a). Increasingly, HCI research has grappled with the challenges of designing user experiences for AI systems particularly due to (1) the uncertainty of AI systems’ capabilities and (2) the complexity of AI systems’ outputs, which may range from simple to adaptively complex (Yang et al., 2020; Dove et al., 2017; Yang et al., 2016). This context of variability and complexity underscores the relevance of DECAI’s approach, which integrates control systems theory to respond to the dynamic nature of sociotechnical systems and their often overlooked feedback loops.
@@ 6. Discussion
Regulatory opportunities. Audits of dark patterns have gained notable regulatory traction, being codified into both EU and US law (European Parliament and the Council of the European Union, 2023b, a; European Commission, 2023a). We believe that our work can aid further research and evidence collection building on this momentum in the AI context (King and Stephan, 2021). Currently, AI-focused legislation such as the EU AI Act only indirectly addresses interface designs through regulations on transparency, human oversight, and AI-driven manipulation (European Commission, 2023b). However, interface designs may allow for the circumvention of such regulation, for example, by allowing for the strategic placement of transparency disclosures to diminish their visibility. Our results also point to how harmful designs can be used to further data collection, a practice central to today’s AI industry (Kak and West, 2023). In that way, interface designs can be pivotal in shaping both current and future regulations aimed at safe and responsible AI practices.